OpenStudy (anonymous):

5 MEDALS help!!!! What is one of the solutions to the following system of equations? y2 + x2 = 65 y + x = 7

OpenStudy (anonymous):

y2 + x2 = 65 y + x = 7 y + x = 7 x = 7 - y Substitute into y^2 + x^2 = 65 y^2 + (7 - y)^2 = 65 y^2 + 49 - 14y + y^2 = 65 2y^2 - 14y - 16 = 0 Divide by 2 y^2 - 7y - 8 = 0 (y + 1)(y - 8) = 0 y = -1 or 8. Substitute into x + y = 7 x = -1 or 8 I hope this helps. :D

OpenStudy (anonymous):

yes thank you so much.

OpenStudy (anonymous):

wouldn't it be -8,1?

OpenStudy (anonymous):

no I don't think so

hero (hero):

@maaryliin you didn't give the person 5 medals